OpenStudy (anonymous):

y varies inversely with twice the value of x. When y = 8, x = 2. Find y when x = 8.

OpenStudy (anonymous):

well x is 1/4 of y so 8 x 4 :)

OpenStudy (anonymous):

Have I failed you Calli?! D:

OpenStudy (callisto):

Hmmm... I don't know y α 1/2x 2xy = k \(2x_1y_1 = 2x_2y_2\) \(2(2)(8) = 2(8)y\) y = ... Now I'm confused...

OpenStudy (anonymous):

so y would be 4? :| I failed you :( lol

OpenStudy (callisto):

No...

OpenStudy (callisto):

Seems I failed myself. I got 2. But.. not sure...

OpenStudy (anonymous):

lol 2 I'm a retard wait 32=16?

OpenStudy (callisto):

2 x 2 x 8 = 2 x 8 x y 32 = 16y y = 32/16 = 2

OpenStudy (anonymous):

y=k/(2x) Put y=8 and x=2, k=32 then y=32/(2x) = 16/x put x=8, y=2

OpenStudy (callisto):

Thanks!!!!

OpenStudy (anonymous):

you are welcome :D